# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 1 Introduction
Rotational motion of the head in humans is perceived through the vestibular system, which is located in the inner ear (Paulin & Hoffman, 2019). For mathematical modeling purposes, this system can be described as a set of three mutually orthogonal, nearly circular canals, known as the semicircular canals (SCCs) in the anatomical literature (Oghalai & Brownell, 2020; Curthoys & Oman, 1987). These canals resemble deformed tori, where the slender regions are filled with a Newtonian fluid called endolymph. The larger region, termed the utricle, houses a gelatinous protein-polysaccharide elastic membrane known as the cupula (Casale et al., 2024), which is innervated by hair cells (cilia); the innervated cilia transmit mechanical deflections of the cupula to the nervous system via the vestibular nerve (Waxman, 2024). In mechanical terms, a change in angular velocity about a given axis drives a fluid flow in that canal, which generates a pressure gradient, deforming the cupula (and hence the innervated hairs) and thus allowing the brain to sense the motion. Specifically, as the walls of the canal rigidly rotate with the head, the fluid in the center is left behind, causing the cupula to deform in the opposite direction to the imposed rotation. The mutually orthogonal structure of the semicircular canals allows the detection of any three dimensional rotation of the head. A schematic diagram of the vestibular system is provided in Figure 1.
@@
A number of mathematical models exist for the vestibular dynamics, both numerical and analytical. On the analytical side, previous models can be classified into two broad categories. The first approach is that of Obrist and co-authors (Obrist, 2008; Grieser et al., 2012; Vega et al., 2008) in which the geometry is idealized to allow a solution to be found under arbitrary forcing of the system, i.e. arbitrary rotational motion. The second approach is that of Rabbitt & Damiano, who maintain a more realistic geometry but require strong assumptions on the form of the forcing to make analytical progress. In particular, in their series of papers (Rabbitt & Damiano, 1992; Damiano & Rabbitt, 1996; Damiano, 1999), Rabbitt & Damiano assume that the forcing is sinusoidal, as might be expected when moving the head up or down (for instance when nodding). Fully numerical investigations of the vestibular system also exist (Chung, 2010; Goyens et al., 2019; Boselli et al., 2009, 2013) — these generally implement a realistic channel geometry but do not model the cupular deformation as a fluid-structure interaction. Instead, they incorporate the effect of cupular deformation via a periodic boundary condition for the flow coupled to a time dependent pressure jump.
In this paper, we present both numerical simulations and an analytical approach for cupular dynamics. Our analytical model is derived from first principles, including explicitly both toroidal fluid flow and the mechanics of the cupula. By exploiting the slenderness of the semicircular canals, and applying a detailed asymptotic analysis, we obtain a reduced model that allows to incorporate both arbitrary geometry and arbitrary forcing, combining the best of previous approaches. We complement this with numerical computations, specifically including fluid-solid couplings. Our combined numerical and analytical approach enables us to validate the reduced analytical model, uncover qualitative features, characterize flow regimes, and investigate detailed features of the flow and cupula deformation.
@@
Moreover, we will explain how the second of these regimes is connected to a symmetry breaking of the flow in the endolymph. As we shall demonstrate in our numerical simulations, presented in Section 2.2, the flow in the endolymph is only axisymmetric (relative to the duct’s center-line or axis) under particular conditions. Despite this observation, many existing models have implicitly assumed a symmetric flow. Examination of our analytical solution enables us to explain exactly when and how symmetry breaking occurs. This feature is interesting in the context of the broader literature on flow through curved pipes. Although the effect of curvature on pipe flow was first discussed by Dean (1928), the plethora of subsequent studies have focused mostly on steady flows (Chupin & Stepanov, 2008; Pedley, 1980; Siggers & Waters, 2005) — there have been many fewer investigations into unsteady fluid phenomena (though see Siggers & Waters, 2008, for an example). We shall show that the essential coupling between the Euler force and the (a priori) unknown pressure gradient can lead to the annihilation of the symmetric leading order velocity — a situation that distinguishes this problem from classical studies of flow in curved pipes. We conclude our study with an investigation of the emergence of vortical flow in the utricle. This feature has been reported previously, but only in numerical simulations (Boselli et al., 2013); our model provides both an analytical understanding and an explicit characterization for when vortical flow will emerge.
@@ 2 Governing equations
Figure 2: Problem setup. (a) Plan view of a semicircular canal showing the spatially-varying canal radius, $a(s)$, and the cupula (shaded in grey), which is situated in the enlarged portion, or utricle. (b) Schematic of the chosen coordinate system. (c) Close up of the region around the cupula, highlighting the cupula’s thickness, $t_{h}$, and its attachment to the canal walls via the ‘crista’ (black region). (The toroidal flow is shown schematically here to allow the zoom in on the cupula.)
We consider a single semicircular canal, as portrayed schematically in Figure 2(a): the endolymph fills a toroidal structure whose centreline forms a circle of radius $R$ and whose radius is small and spatially-varying, denoted $\hat{a}(\hat{s})\ll R$, where $\hat{s}$ is an arc length parameter along the centreline. The canal is subjected to a rotation defined by angular velocity $\hat{\Omega}(\hat{t})$ around the centre of the toroid with rotation axis normal to the plane of the centreline. The endolymph is assumed to be an incompressible Newtonian fluid of dynamic viscosity $\mu$ and density $\rho$. The elastic, gel-like cupula occupies a thin region (shown in grey in fig. 2) and has density $\rho_{s}$, Young’s modulus $E$, thickness $t_{h}$ and Poisson ratio $\nu_{s}$. The cupula is assumed to occupy the entire cross-section of the canal, as can be seen in Figure 2(c), where the solid cupula is shaded in yellow and the liquid endolymph is shaded in blue. (The region shaded in black represents a structure called the crista which attaches the cupula to the canal wall. However, the azimuthal size of the crista will not appear directly in this work.)
@@ 2.1 Equations for the bulk
The Navier–Stokes equations for the dimensional fluid velocity $\hat{\mathbf{u}}$ and pressure $\hat{p}$ in the co-rotating frame are given by (Landau & Lifshitz, 1987)
@@
$\frac{D\mathbf{\hat{u}}}{D\hat{t}}+\frac{\partial\mathbf{\hat{\Omega}}}{\partial\hat{t}}\times\mathbf{\hat{x}}+2\mathbf{\hat{\Omega}}\times\mathbf{\hat{u}}+\mathbf{\hat{\Omega}}\times(\mathbf{\hat{\Omega}}\times\mathbf{\hat{x}})$ $=-\boldsymbol{\nabla}\hat{p}+\mu\boldsymbol{\nabla}^{2}\mathbf{\hat{u}}.$ (1b)
Here, the first of the extra terms on the left hand side is the Euler force, the second term is the Coriolis force and the final additional term is the centrifugal force, each due to the imposed rotation. The fluid is assumed to satisfy the no-slip condition at the edges of the walls, so that $\mathbf{\hat{u}}(\hat{r}=\hat{a}(\hat{s}))=\mathbf{0}$ for $s\in\,(0,2\pi R)$. Motivated by the small strains in the cupula (Selva et al., 2009), it is modeled as a linearly elastic material, satisfying the steady Navier equations:
$\boldsymbol{\nabla}\cdot\boldsymbol{\hat{\tau}}=\mathbf{0},$ (2)
alongside a linear Hookean constitutive law relating stress $\boldsymbol{\hat{\tau}}$ and the strains. Finally, we model the fluid structure interaction at the cupula-endolymph boundary in the usual way, imposing continuity of velocity and stress (Gkanis & Kumar, 2006).
@@ 2.2 Numerical simulations
The system of equations (1)-(2) was simulated in COMSOL for different values of the Young’s modulus of the cupula and with a Poisson ratio $\nu_{s}=0.3$. We impose a simple sinusoidal forcing, given by $\hat{\Omega}(\hat{t})=\Omega_{0}\sin(2\pi t/t_{\rm typ})$, with $\Omega_{0}=1\mathrm{rad/s}$ and $t_{\rm typ}=1\mathrm{~{}s}$. The geometrical parameters are $a=1.6\times 10^{-4}$ m, $R=3.2\times 10^{-3}$ m and $t_{h}=0.8\times 10^{-4}$ m (Daocai et al., 2014).
@@
The equations were solved for a solid cupula of finite size, i.e. with no thin cupula assumption. Further details on the numerical simulations are available in Appendix A. This includes deformation profiles of the solid cupula (Figure 12). Moreover, Figure 12 shows that the magnitude of cupula deformation is inversely proportional to Young’s modulus.
Flow profiles on either side of the cupula produced by these numerical simulations are shown in fig. 3. Here a top view of the mid-plane of the flow around the canal is plotted, with colour indicating the speed of the flow, normalized by the maximum speed throughout the flow and with fast regions coloured red, stagnant regions coloured blue.
The cupula appears as the central region and is shown in its deformed configuration. This deflection is imperceptible for all except the $E=10^{2}\mathrm{~{}Pa}$ case and so the relative magnitude of the cupula deformation is indicated by the grey scale colouring, which shows that it is maximum at the centre.
@@
In the panels of fig. 3, the Young’s modulus, $E$, of the cupula increases from left to right. As should be expected, the magnitude of the deflection decreases as the material becomes stiffer. This is confirmed quantitatively in Figure 12 of Appendix A. Surprisingly, however, we also observe a significant change in flow behaviour: for small values of $E$, the flow is axially symmetric about the centerline of the tube but as $E$ increases the flow transitions, losing its axisymmetry and exhibiting vortical structures. It is worth noting that these plots use values of the Young’s modulus that cover a typical physiological range for soft biological tissues (Goriely, 2017). Notwithstanding our remark in the introduction that even lower values have been suggested via indirect methods, a flow transition with physiologically relevant values of $E$ suggests that asymmetric flow may be present in a physiological vestibular system. If so, this would contradict the assumption of axial symmetry that is typical in previous models (Obrist, 2008; Rabbitt & Damiano, 1992), and raises interesting questions about what impact such asymmetry might have on the mechanics of balance and rotational sensing. To study this behaviour further, we thus turn now to a theoretical analysis of the governing system.
Figure 3: Velocity fields in the cupula as computed using COMSOL simulations. Results are shown for a range of cupula stiffnesses. Colour represents the relative magnitude of the fluid speed, with red denoting regions in which the flow is fast and blue representing stagnant regions; streamlines are represented by solid black curves. As the stiffness of the cupula increases, a symmetry-breaking of the flow occurs. In particular, for values of the Young’s modulus $E>10^{3}$ the flow is usually not axially-symmetric. Here $\hat{\Omega}(\hat{t})=\sin\left(2\pi\hat{t}/\mathcal{T}\right)$ and the snapshots are taken at $\hat{t}=0.25$ s and the period is $\mathcal{T}=1$ sec.
@@ 2.3 Theoretical model
To capture the geometry of the semicircular canal, we introduce toroidal coordinates $(\hat{r},\theta,\hat{s})$, in which $\hat{s}\,\in\,(0,2\pi R)$ is the arc-length along the centreline of the torus. A sketch of this coordinate system is provided in Figure 2(b). Cartesian coordinates are related to toroidal coordinates via (Pedley, 1980):
$\begin{cases}\hat{x}=&(R+\hat{r}\cos{\theta})\cos\frac{\hat{s}}{R},\\ \hat{y}=&(R+\hat{r}\cos\theta)\sin\frac{\hat{s}}{R},\\ \hat{z}=&-\hat{r}\sin\theta.\end{cases}$ (3)
@@
$\begin{split}\rho&\left(\frac{\partial\hat{u}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{u}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\hat{v}^{2}}{\hat{r}}-\frac{\hat{w}^{2}}{h}\frac{\cos\theta}{R}-\hat{\Omega}\hat{w}\cos\theta-\hat{\Omega}^{2}Rh\cos\theta\right)\\ &=-\frac{\partial\hat{p}}{\partial\hat{r}}+\frac{\mu}{\hat{r}h}\left[\frac{\hat{r}}{h}\frac{\partial}{\partial\hat{s}}\left(\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\partial}{\partial\hat{r}}(h\hat{w})\right)-\frac{1}{\hat{r}}\frac{\partial}{\partial\theta}\left(\frac{h}{\hat{r}}\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{h}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}\right)\right],\end{split}$ (5a)
@@
$\begin{split}\rho&\left(\frac{\partial\hat{v}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{v}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{v}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{v}}{\partial\hat{s}}+\frac{\hat{u}\hat{v}}{\hat{r}}+\frac{\hat{w}^{2}}{Rh}\sin\theta+\hat{\Omega}\hat{w}\sin\theta+\hat{\Omega}^{2}Rh\sin\theta\right)\\ &=-\frac{1}{\hat{r}}\frac{\partial\hat{p}}{\partial\theta}+\frac{\mu}{h}\frac{\partial}{\partial\hat{r}}\left[\frac{h}{\hat{r}}\left(\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{\partial\hat{u}}{\partial\theta}\right)\right]-\frac{\mu}{\hat{r}h^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial\hat{w}}{\partial\hat{s}}\right)-\hat{r}\frac{\partial^{2}\hat{v}}{\partial\hat{s}^{2}}\right],\end{split}$ (5b)
$\begin{split}\rho&\left(\frac{\partial\hat{w}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{w}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{w}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{w}}{\partial\hat{s}}+\frac{\hat{u}\hat{w}}{Rh}\cos\theta-\frac{\hat{v}\hat{w}}{Rh}\sin\theta\right.\\ &\left.+\frac{d\hat{\Omega}}{d\hat{t}}(R+\hat{r}\cos\theta)+\hat{\Omega}\hat{u}\cos\theta-\hat{\Omega}\hat{v}\sin\theta\right)\\ &=-\frac{1}{h}\frac{\partial\hat{p}}{\partial\hat{s}}+\frac{\mu}{\hat{r}^{2}}\frac{\partial}{\partial\theta}\left[\frac{1}{h}\frac{\partial}{\partial\theta}(h\hat{w})-\frac{\hat{r}}{h}\frac{\partial\hat{v}}{\partial\hat{s}}\right]-\frac{\mu}{\hat{r}}\frac{\partial}{\partial\hat{r}}\left[\frac{\hat{r}}{h}\left(\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\partial}{\partial\hat{r}}(h\hat{w})\right)\right].\end{split}$ (5c)
Noting that the cupula is approximately thin (with an aspect ratio of around 0.25), we model it as a circular plate. We locate the cupula at arc length position $\hat{s}=0$, and define its deflection by the function $\hat{\eta}(\hat{r},\theta,\hat{t})$. As indicated above, our numerical simulations have shown that the deformation of the cupula is small compared to the tube radius, suggesting that strains are small and suggesting it is sufficient to use a linear equation. This small strain assumption will be confirmed in Section 2.4 through a scaling argument. Motivated by this linearity and the relatively slender geometry of the cupula, we further assume that the cupular deflection satisfies the linear plate equation, which in the rotating frame reads (see Howell et al., 2008)
$B\nabla^{4}\hat{\eta}-T\nabla^{2}\hat{\eta}+\rho_{s}t_{h}\left(\frac{\partial^{2}\hat{\eta}}{\partial\hat{t}^{2}}+Rh\frac{\mathrm{d}\hat{\Omega}}{\mathrm{d}\hat{t}}\right)=\Delta\hat{p}=\hat{p}(\hat{r},\theta,\hat{s}=2\pi R,\hat{t})-\hat{p}(\hat{r},\theta,\hat{s}=0,\hat{t}).$ (6)
where $B=Et_{h}^{3}/[12(1-\nu_{s}^{2})]$ is the plate’s bending stiffness and $T$ is the applied tension in the plate (which we assume is uniform and constant since deflections remain small compared to $t_{h}$). In this regime, the forcing for the cupular displacement is given by the pressure jump across the two sides of the cupula.
@@
$\frac{\partial\hat{\eta}}{\partial\hat{t}}\boldsymbol{e}_{s}=\mathbf{\hat{u}}(\hat{r},\hat{s}=0,\hat{t})=\mathbf{\hat{u}}(\hat{r},\hat{s}=2\pi R,\hat{t}).$ (7)
We also require boundary conditions for the plate equation (6). The precise attachment between the cupula and the utricle is an open area of research and the conditions to be satisfied are not immediately clear. We opt to implement a straightforward choice suggested by Rabbitt & Damiano (1992), which is to apply clamped boundary conditions at the edge of the cupula, i.e.
$\hat{\eta}(\hat{r}=\hat{a}(0))=\left.\frac{\partial\hat{\eta}}{\partial\hat{r}}\right|_{\hat{r}=\hat{a}(0)}=0.$ (8)
We also require initial conditions for the velocity, $\mathbf{\hat{u}}(\mathbf{\hat{x}},\hat{t}=0)$ and the cupular deflection $\hat{\eta}(\hat{r},\theta,\hat{t}=0)$. Unless otherwise stated we assume that the system is initially at rest and is undeformed.

### v2

@@ 1 Introduction
Rotational motion of the head in humans is perceived through the vestibular system, which is located in the inner ear (Paulin & Hoffman, 2019). For mathematical modeling purposes, this system can be described as a set of three mutually orthogonal, nearly circular canals, known as the semicircular canals (SCCs) in the anatomical literature (Oghalai & Brownell, 2020; Curthoys & Oman, 1987). These canals resemble deformed tori, where the slender regions are filled with a Newtonian fluid called endolymph. The larger region is comprised of two cavities, the utricle and the ampulla. The latter houses a gelatinous protein-polysaccharide elastic membrane known as the cupula (Casale et al., 2024), which is innervated by hair cells (cilia); the innervated cilia transmit mechanical deflections of the cupula to the nervous system via the vestibular nerve (Waxman, 2024). In mechanical terms, a change in angular velocity about a given axis drives a fluid flow in that canal, which generates a pressure gradient, deforming the cupula (and hence the innervated hairs) and thus allowing the brain to sense the motion. Specifically, as the walls of the canal rigidly rotate with the head, the fluid in the center is left behind, causing the cupula to deform in the opposite direction to the imposed rotation. The mutually orthogonal structure of the semicircular canals allows the detection of any three dimensional rotation of the head. A schematic diagram of the vestibular system is provided in figure 1.
@@
A number of mathematical models exist for the vestibular dynamics, both numerical and analytical. On the analytical side, beyond the early phenomenological oscillator models from the 1930’s (Steinhausen, 1933), recent models can be classified into two broad categories. The first approach is that of Obrist and co-authors (Obrist, 2008; Vega et al., 2008; Buskirk et al., 1976) in which the geometry is idealized to allow a solution to be found under arbitrary forcing of the system, i.e. arbitrary rotational motion. The second approach is that of Rabbitt & Damiano, who maintain a more realistic geometry but require strong assumptions on the form of the forcing to make analytical progress. In particular, in their series of papers (Rabbitt & Damiano, 1992; Damiano & Rabbitt, 1996; Damiano, 1999), Rabbitt & Damiano assume that the forcing is sinusoidal, as might be expected when tilting the head up or down (for instance when nodding). Fully numerical investigations of the vestibular system also exist (Boselli et al., 2009; Grieser et al., 2012; Boselli et al., 2013) — these generally implement a realistic channel geometry but do not model the cupular deformation as a fluid-structure interaction. Instead, they incorporate the effect of cupular deformation via a periodic boundary condition for the flow coupled to a time dependent pressure jump. Recent studies have also modelled the fluid–structure interaction (Chung et al., 2010; Goyens et al., 2019; Kassemi et al., 2005; Wu et al., 2011); however, the high computational cost of such simulations often restricts them to a single manoeuvre and a single set of parameters, rather than enabling a broader exploration of parameter space to uncover the flow and deformation regimes of the vestibular system.
In this paper, we present both numerical simulations and an analytical approach for cupular dynamics. Our analytical model is derived from first principles, including explicitly both toroidal fluid flow and the mechanics of the cupula. By exploiting the slenderness of the semicircular canals, and applying a detailed asymptotic analysis, we obtain a reduced model that allows us to incorporate both arbitrary geometry and arbitrary forcing, combining the best of previous approaches. Moreover, unlike previous work (Rabbitt & Damiano, 1992; Obrist, 2008) we model the cupula as a full three-dimensional elastic solid, without making a small thickness assumption. We complement this with numerical computations, specifically including fluid-solid couplings. Our combined numerical and analytical approach enables us to validate the reduced analytical model, uncover a number of novel features, including characterizing flow regimes and identifying regions of parameter space with distinct system response.
@@
Moreover, we will explain how the second of these regimes is connected to a symmetry breaking of the flow in the endolymph. As we shall demonstrate in our numerical simulations, presented in Section 2.2, the flow in the endolymph is only axisymmetric (relative to the duct’s center-line or axis) under particular conditions. Despite this observation, many existing models have implicitly assumed a radially symmetric flow. Examination of our analytical solution enables us to explain exactly when and how symmetry breaking occurs. This feature is interesting in the context of the broader literature on flow through curved pipes. Although the effect of curvature on pipe flow was first discussed by Dean (1928), the plethora of subsequent studies have focused mostly on steady flows (Chupin & Stepanov, 2008; Pedley, 1980; Siggers & Waters, 2005) — there have been many fewer investigations into unsteady fluid phenomena (though see Siggers & Waters, 2008, for an example). We shall show that the essential coupling between the Euler force and the (a priori) unknown pressure gradient can lead to the annihilation of the symmetric leading order velocity — a situation that distinguishes this problem from classical studies of flow in curved pipes. We conclude our study with an investigation of the emergence of vortical flow in the utricle. This feature has been reported previously, but only in numerical simulations (Grieser et al., 2012; Boselli et al., 2013); our model provides both an analytical understanding and an explicit characterization for when vortical flow will emerge.
@@ 2 Governing equations
[…]
@@ 2.1 Equations for the bulk
The Navier–Stokes equations for the dimensional fluid velocity $\hat{\mathbf{u}}$ and modified pressure $\hat{p}$ in the co-rotating frame are given by (Landau & Lifshitz, 1987)
@@
[…]
Here, the first of the extra terms on the left hand side is the Euler force, the second term is the Coriolis force and the final additional term is the centrifugal force, each due to the imposed rotation. The pressure $\hat{p}$ is a modified pressure in the sense that it incorporates the linear fictitious force associated with the linear acceleration of the SCC (Buskirk et al., 1976). The fluid is assumed to satisfy the no-slip condition at the edges of the walls, so that $\mathbf{\hat{u}}(\hat{r}=\hat{a}(\hat{s}))=\mathbf{0}$ for $\hat{s}\in\,(0,2\pi R)$. Motivated by the small strains in the cupula (Selva et al., 2009), it is modeled as a linearly elastic material, satisfying the steady Navier equations:
$\rho_{s}\left(\frac{\partial^{2}\boldsymbol{\hat{u}}_{s}}{\partial t^{2}}+\frac{\partial\mathbf{\hat{\Omega}}}{\partial\hat{t}}\times\mathbf{\hat{x}}+2\mathbf{\hat{\Omega}}\times\frac{\partial\boldsymbol{\hat{u}}_{s}}{\partial\hat{t}}+\mathbf{\hat{\Omega}}\times(\mathbf{\hat{\Omega}}\times\mathbf{\hat{x}})\right)=\boldsymbol{\nabla}\cdot\boldsymbol{\hat{\tau}}$, (2a)
$\boldsymbol{\hat{\tau}}=2\mu_{s}\boldsymbol{\hat{E}}+\lambda_{s}\mathrm{tr}(\boldsymbol{\hat{E}})\mathbf{1},\quad 2\boldsymbol{\hat{E}}=\boldsymbol{\nabla}\boldsymbol{\hat{u}}_{s}+(\boldsymbol{\nabla}\boldsymbol{\hat{u}}_{s})^{\intercal},$ (2b)
alongside a linear Hookean constitutive law relating stress $\boldsymbol{\hat{\tau}}$ and the cupular displacement field $\boldsymbol{\hat{u}}_{s}(\mathbf{x},t)$. The Lame constants are $\mu_{s}=E/(2(1+\nu_{s}))$ and $\lambda_{s}=\nu_{s}E/((1+\nu_{s})(1-2\nu_{s}))$ (Bower, 2009). Finally, we model the fluid structure interaction at the cupula-endolymph boundary in the usual way, imposing continuity of velocity and stress (Gkanis & Kumar, 2006).
@@ 2.2 Numerical simulations
The system of equations (1)-(2) was simulated in COMSOL for different values of the Young’s modulus of the cupula and with a Poisson ratio $\nu_{s}=0.48$. We impose a simple sinusoidal forcing, given by $\hat{\Omega}(\hat{t})=\Omega_{0}\sin(2\pi\hat{t}/\mathcal{T})$, with $\Omega_{0}=1\,\mathrm{rad/s}$ and $\mathcal{T}=1\mathrm{~s}$. The geometrical parameters are $a=1.6\times 10^{-4}$ m, $R=3.2\times 10^{-3}$ m and $t_{h}=0.8\times 10^{-4}$ m (Daocai et al., 2014).
@@
The equations were solved for a solid cupula of finite size, i.e. with no thin cupula assumption. Further details on the numerical simulations are available in Appendix B.1. This includes deformation profiles of the solid cupula (Figure 13). Moreover, Figure 13 shows that the magnitude of cupula deformation is inversely proportional to Young’s modulus.
Flow profiles on either side of the cupula produced by these numerical simulations are shown in figure 3. Here a top view of the mid-plane of the flow around the canal is plotted, with colour indicating the speed of the flow, normalized by the maximum speed throughout the flow and with fast regions coloured red, stagnant regions coloured blue. The cupula appears as the central region and is shown in its deformed configuration. This deflection is imperceptible for all except the $E=10^{2}\mathrm{~Pa}$ case and so the relative magnitude of the cupula deformation is indicated by the grey scale colouring, which shows that it is maximum at the centre. Figure 3 shows the velocity field at time $\hat{t}=0.25$ s, a stage at which transients from the initial condition remain significant. However, it is not necessary to wait for these transients to decay completely to observe the emergence of a distinctly asymmetric velocity profile.
@@
In the panels of figure 3, the Young’s modulus, $E$, of the cupula increases from left to right. As should be expected, the magnitude of the deflection decreases as the material becomes stiffer. This is confirmed quantitatively in Figure 13 of Appendix B.1. Surprisingly, however, we also observe a significant change in flow behaviour: for small values of $E$, the flow is axially symmetric about the centerline of the tube but as $E$ increases the flow transitions, losing its axisymmetry and exhibiting vortical structures. It is worth noting that these plots use values of the Young’s modulus that cover a typical physiological range for soft biological tissues (Goriely, 2017). Notwithstanding our remark in the introduction that even lower values have been suggested via indirect methods, a flow transition with physiologically relevant values of $E$ suggests that asymmetric flow may be present in a physiological vestibular system. If so, this would contradict the assumption of axial symmetry that is typical in previous models (Obrist, 2008; Rabbitt & Damiano, 1992), and raises interesting questions about what impact such asymmetry might have on the mechanics of balance and rotational sensing. To study this behaviour further, we thus turn now to a theoretical analysis of the governing system.
[…]
@@ 2.3 Theoretical model
[…]
$\begin{cases}\hat{x}=&(R+\hat{r}\cos{\theta})\cos(\frac{\hat{s}}{R}),\\ \hat{y}=&(R+\hat{r}\cos\theta)\sin(\frac{\hat{s}}{R}),\\ \hat{z}=&-\hat{r}\sin\theta.\end{cases}$ (3)
@@
$\begin{split}\rho&\left(\frac{\partial\hat{u}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{u}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\hat{v}^{2}}{\hat{r}}-\frac{\hat{w}^{2}}{h}\frac{\cos\theta}{R}-2\hat{\Omega}\hat{w}\cos\theta-\hat{\Omega}^{2}Rh\cos\theta\right)\\ &=-\frac{\partial\hat{p}}{\partial\hat{r}}+\frac{\mu}{\hat{r}h}\left[\frac{\hat{r}}{h}\frac{\partial}{\partial\hat{s}}\left(\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\partial}{\partial\hat{r}}(h\hat{w})\right)-\frac{\partial}{\partial\theta}\left(\frac{h}{\hat{r}}\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{h}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}\right)\right],\end{split}$ (5a)
@@
$\begin{split}\rho&\left(\frac{\partial\hat{v}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{v}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{v}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{v}}{\partial\hat{s}}+\frac{\hat{u}\hat{v}}{\hat{r}}+\frac{\hat{w}^{2}}{Rh}\sin\theta+2\hat{\Omega}\hat{w}\sin\theta+\hat{\Omega}^{2}Rh\sin\theta\right)\\ &=-\frac{1}{\hat{r}}\frac{\partial\hat{p}}{\partial\theta}+\frac{\mu}{h}\frac{\partial}{\partial\hat{r}}\left[\frac{h}{\hat{r}}\left(\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{\partial\hat{u}}{\partial\theta}\right)\right]-\frac{\mu}{\hat{r}h^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial\hat{w}}{\partial\hat{s}}\right)-\hat{r}\frac{\partial^{2}\hat{v}}{\partial\hat{s}^{2}}\right],\end{split}$ (5b)
[…]
We locate the cupula at arc length position $\hat{s}=0$. As indicated above, our numerical simulations have shown that the deformation of the cupula is small compared to the tube radius, suggesting that strains are small and thus that it is sufficient to use a linear equation. This small strain assumption will be confirmed in Section 2.4 through a scaling argument. We write the equations for the solid deformation of the cupula in cylindrical coordinates, as the cupula is thin enough that the curvature of the SCC is not important. The cupula is thus modelled as a solid cylinder $0\leq\hat{r}\leq\hat{a}(0)$, $0\leq\theta\leq 2\pi$, $t_{h}/2\leq\hat{z}\leq t_{h}/2$, with the centre of mass of the cupula ($\hat{r}=\theta=\hat{z}=0$) located at position $\hat{r}=\theta=\hat{s}=0$ (or equivalently $\hat{s}=2\pi R$) in terms of the toroidal coordinates of the fluid problem (see figure 2(a)). The thickness of the cupula $t_{h}$ is not assumed to be small when compared to its radius $\hat{a}(0)$. The equations for the solid deformation in component form (2) are given in §B.2.
@@
[…]
We also require boundary conditions for the solid problem  (2). The precise attachment between the cupula and the utricle is an open area of research and the conditions to be satisfied are not immediately clear. We opt to implement a straightforward choice motivated by Rabbitt & Damiano (1992) who model the cupula as a clamped plate, but we extend this to include thickness effects. In this regime, the forcing for the cupular displacement is given by the pressure jump across the two sides of the cupula. In particular, at the flat faces of the cylinder we impose the following conditions on the stress tensor
$\hat{\tau}_{zz}=\pm\frac{1}{2}\left[\hat{p}(\hat{r},\theta,\hat{s}=2\pi R-t_{h}/2,\hat{t})-\hat{p}(\hat{r},\theta,\hat{s}=t_{h}/2,\hat{t})\right]\equiv\pm\frac{\Delta\hat{p}}{2},\quad$ $\text{at }\hat{z}=\pm t_{h}/2,$ (7a)
$\hat{\tau}_{rz}=0,\quad\hat{\tau}_{\theta z}=0,\quad$ $\text{at }\hat{z}=\pm t_{h}/2.$ (7b)
On the curved face we impose zero displacement, $\boldsymbol{\hat{u}}_{s}(\hat{r}=\hat{a}(0))=\boldsymbol{0}$. We also require initial conditions for the velocity, $\mathbf{\hat{u}}(\mathbf{\hat{x}},\hat{t}=0)$ and the cupular deflection $\boldsymbol{\hat{u}_{s}}(\hat{t}=0)$. Unless otherwise stated we assume that the system is initially at rest and is undeformed.